# Linear and Sublinear Diversities

### v2

@@ 2.2 Properties of linear and sublinear diversities
Let $\delta$ be a function on finite subsets of $\mathbb{R}^{k}$ which satisfies (D1), monotonicity (D3) and sublinearity (D6).
@@
    $\delta$ is translation invariant: $\delta(A+x)=\delta(A)$ for all finite $A\subseteq\mathbb{R}^{k}$ and $x\in\mathbb{R}$.
@@
    The map $N:\mathbb{R}^{k}\rightarrow\mathbb{R}$ given by $N(x)=\delta(\{0,x\})$ is a norm on $\mathbb{R}^{k}$.
@@
If $\delta$ satisfies (D1) rather than (D1) then 1-5 still hold except that $(\mathbb{R}^{k},\delta)$ is a semidiversity and $N$ is a seminorm.
@@
    As $\delta$ is monotonic and $\delta(\emptyset)=0$, $\delta$ is non-negative, and by part 1. $\delta$ is translation invariant. We show that $(X,\delta)$ satisfies (D4). Suppose that $x\in A\cap B$. Then $0\in(A-x)\cap(B-x)$ and so $(A-x)\cup(B-x)\subseteq(A-x)+(B-x)$ and
@@
    Proposition 2.2b in [3].
@@
    By (D5) we have $N(x+y)=\delta(\{0,x+y\})\leq\delta(\{0,x\})+\delta(\{0,y\})=N(x)+N(y)$. If $\lambda>0$ then $N(\lambda x)=\delta(\{0,\lambda x\})=\lambda\delta(0,x)=|\lambda|N(x)$, while if $\lambda<0$ we have
@@
    Also, $N(0)=\delta(\{0\})=0$ if and only if $\delta(\{x,0\})=0$ if and only if $x=0$.
  5. 5.
    This follows from sublinearity and the following observation; see the proof of [1, Theorem 4.1]. By sublinearity we may assume $\sum_{a\in A}a=0$. So for each $a\in A$
    $-\frac{1}{|A|-1}a=\frac{1}{|A|-1}\sum_{a^{\prime}\neq a}a^{\prime}\in\mathrm{conv}(A\setminus\{a\}).$
    We also have $a\in\mathrm{conv}(A\setminus a^{\prime})$ for $a\neq a^{\prime}$. So for all $a\in A$
    $\frac{(|A|-2)|A|}{|A|-1}a=(|A|-1)a-\frac{1}{|A|-1}a\in\sum_{b\in A}\mathrm{conv}(A\setminus\{b\}).$
    This gives
    $A\subseteq\frac{|A|-1}{|A|(|A|-2)}\sum_{a\in A}\mathrm{conv}(A\setminus\{a\})$
    and applying sublinearity gives the result.
It is now straightforward to show that the zonotope diversity introduced above is in fact a sublinear diversity.
@@
Recall that $\delta_{z}(A)$ is the shortest length of a zonotope containing $A$. The function $\delta_{z}(A)$ is clearly monotonic, vanishes when $|A|\leq 1$, and is strictly positive when $|A|>1$. Given finite $A,B$, let $Z_{A}$ and $Z_{B}$ the the minimum length zonotopes containing $A$ and $B$ respectively. Then $Z_{A}+Z_{B}$ is a zonotope containing $A+B$ with length $\ell(Z_{A})+\ell(Z_{B})$. By Proposition 1 part 2, $(\mathbb{R}^{k},\delta_{z})$ is a sublinear diversity. ∎
The zonotope diversity is not linear: let $A=\{(0,0),(1,0),(0,1)\}$ and $B=-A$. Then $\delta_{z}(A)=\delta_{z}(B)=2$ but $\delta_{z}(A+B)=2+\sqrt{2}$.
In a semidiversity, (D1) is replaced by (D1), and sets with more than one element can have diversity zero. When the semidiversity is sublinear, the sets with zero diversity are highly structured. Define the null set of a semidiversity $(\mathbb{R}^{k},\delta)$ to be the set
$\mathrm{null}(\delta)=\Big{\{}x:\delta(\{0,x\})=0\Big{\}}$
and $\mathrm{null}(\delta)^{\perp}=\{x\in\mathbb{R}^{k}:x\cdot y=0\mbox{ for all }y\in\mathrm{null}(\delta)\}$.
Proposition 3.
Let $(\mathbb{R}^{k},\delta)$ be a sublinear semidiversity.
  1. 1.
    $\mathrm{null}(\delta)$ is a linear subspace of $\mathbb{R}^{k}$
  2. 2.
    $\delta$ restricted to $\mathrm{null}(\delta)^{\perp}$ is a diversity
  3. 3.
    If $P$ is the projection operator for $\mathrm{null}(\delta)^{\perp}$ then $\delta(A)=\delta(PA)$ for all finite $A\subseteq\mathbb{R}^{k}$.
Proof.
  1. 1.
    For $x,y\in\mathrm{null}(\delta)$ and $\alpha>0$ we have $\delta(\{0,x+y\})\leq\delta(\{0,x\})+\delta(\{0,y\})=0$ and $\delta(\{0,\alpha x\})=\alpha\delta(\{0,x\})=0$ so $x+y\in\mathrm{null}(\delta)$ and $\alpha x\in\mathrm{null}(\delta)$. By translation invariance, $\delta(\{0,-x\})=\delta(\{x,0\})=0$ and $-x\in\mathrm{null}(\delta)$.
  2. 2.
    Suppose $x,y\in\mathrm{null}(\delta)^{\perp}$ and $\delta(\{x,y\})=0$. We have that $x-y\in\mathrm{null}(\delta)^{\perp}$ by part 1. By translation invariance $\delta(\{0,x-y\})=\delta(\{x,y\})=0$ which implies $x-y\in\mathrm{null}(\delta)$. Hence $x-y$ is both in a subspace and its orthogonal complement, and so $x=y$.
  3. 3.
    For all finite $A\subseteq\mathbb{R}^{k}$ we have $A\subseteq PA+B$ and $PA\subseteq A+C$ for some $B,C\subset\mathrm{null}(\delta)$. We have $\delta(B)=0$ since
    $0\leq\delta(B)\leq\sum_{b\in B}\delta(\{0,b\})=0,$
    and, likewise, $\delta(C)=0$. By sublinearity, $\delta(A)=\delta(PA)$.
[…]
@@
For bounded $K\subseteq\mathbb{R}^{k}$ define
@@
    For all finite $A\subseteq\mathbb{R}^{k}$ we have
@@
    By equivalency of norms on $\mathbb{R}^{k}$ we have that $K$ is bounded with respect to metric $d$ if and only if it is bounded with respect to the induced metric of $\delta$. Let $V$ be the set of vertices of some polytope (e.g. a cube) containing $K$. For all finite $A\subseteq K$ we have by monotonicity and part 2 that
    $\delta(A)\leq\delta(A\cup V)=\delta(V)$
    so that $\delta^{*}(K)\leq\delta(V)<\infty.$
@@
    Let $A$ be a finite subset of $\mathbb{R}^{k}$ and let $K=\mathrm{conv}(A)$. For any $A^{\prime}\subseteq K$ we have $\mathrm{conv}(A\cup A^{\prime})=\mathrm{conv}(A)$ so $\delta(A^{\prime})\leq\delta(A\cup A^{\prime})=\delta(A)$ by Proposition 1 (ii). Hence
    $\delta(A)\leq\sup\{\delta(A^{\prime}):\mbox{ finite }A^{\prime}\subseteq K)\}\leq\delta(A).$
@@
    Let $A$ be a finite subset of $K$ such that $\delta(A)>\delta^{*}(K)-\epsilon$. As $\lambda A\subseteq\lambda K$ we have
    $\lambda(\delta^{*}(K)-\epsilon)<\lambda\delta(A)=\delta(\lambda A)\leq\delta^{*}(\lambda K).$
    Hence $\delta^{*}(\lambda K)\geq\lambda\delta^{*}(K)$ from which equality follows by symmetry.
@@
    By Lemma 3.1.1 of [24] we have that if $K,L,K\cup L$ and $K\cap L$ are nonempty compact convex subsets then
@@
    Suppose that $K,L$ are bounded nonempty subsets satisfying $d_{H}(K,L)=\lambda$. For any $\epsilon>0$ there is a finite $A\subseteq K$ such that $\delta(A)\leq\delta^{*}(K)<\delta(A)+\epsilon$. We also have $A\subseteq K\subseteq L+\lambda\mathcal{B}$ so there is finite $B\subseteq L$ and $C\subseteq\mathcal{B}$ such that $A\subseteq B+\lambda C$. Hence
@@
Bryant et al. [3] also describe an extension of Minkowski diversities from finite sets to bounded sets. They define $\widetilde{\delta}(P)=\delta(\mathrm{vert}(P))$ for any polytope with vertex set $\mathrm{vert}(P)$, and extend that to general bounded convex sets $K$ by defining $\widetilde{\delta}(K)=\lim_{n\rightarrow\infty}\widetilde{\delta}(P_{n})$ for any sequence of polytopes $P_{1},P_{2},\ldots$ converging to $K$. Proposition 4 part 2. gives that $\delta^{*}(P)=\widetilde{\delta}(P)$ for any polytope, while from Proposition 4 part 5 we have that $\delta^{*}(K)=\widetilde{\delta}(K)$. Hence $\delta^{*}$ coincides with $\widetilde{\delta}$ for Minkowski diversities.

### v3

@@ 2.2 Properties of linear and sublinear diversities
Let $\delta$ be a function on finite subsets of $\mathbb{R}^{k}$ which satisfies (D1), monotonicity (D3) and sublinearity (D5). Then the following hold:
@@
    $\delta$ is translation invariant: $\delta(A+x)=\delta(A)$ for all finite $A\subseteq\mathbb{R}^{k}$ and $x\in\mathbb{R}^{k}$.
@@
    If $A\subseteq\mathrm{conv}(B)$ then $\delta(A)\leq\delta(B)$.
@@
    The map $N:\mathbb{R}^{k}\rightarrow\mathbb{R}$ given by $N(x)=\delta(\{0,x\})$ is a norm on $\mathbb{R}^{k}$ such that $\delta(\{x,y\})=N(x-y)$ for all $x,y\in\mathbb{R}^{k}$.
  6. 6.
@@
If $\delta$ satisfies (D1) rather than (D1) then 1. through 6. still hold except that $(\mathbb{R}^{k},\delta)$ is a semidiversity and $N$ is a seminorm.
@@
    As $\delta$ is monotonic and $\delta(\emptyset)=0$, $\delta$ is non-negative, and by part 1., $\delta$ is translation invariant. We show that $(\mathbb{R}^{k},\delta)$ satisfies (D4). Suppose that $x\in A\cap B$. Then $0\in(A-x)\cap(B-x)$ and so $(A-x)\cup(B-x)\subseteq(A-x)+(B-x)$ and
@@
    By part 2., $(X,\delta)$ is a sublinear diversity. Then by (bryant2023diversities, , Prop. 2.2b), if $A,B$ are finite sets with $\mathrm{conv}(A)=\mathrm{conv}(B)$ then $\delta(A)=\delta(B)$.
@@
    We have $\mathrm{conv}(A\cup B)=\mathrm{conv}(B)$ so by 3., $\delta(B)=\delta(A\cup B)\geq\delta(A)$.
  5. 5.
    By (D5) we have $N(x+y)=\delta(\{0,x+y\})\leq\delta(\{0,x\})+\delta(\{0,y\})=N(x)+N(y)$. If $\lambda\geq 0$ then $N(\lambda x)=\delta(\{0,\lambda x\})=\lambda\delta(\{0,x\})=|\lambda|N(x)$, while if $\lambda<0$ we have
@@
    Also, since $N(x)=\delta(\{0,x\})$, $N(x)=0$ if and only if $\delta(\{x,0\})=0$ if and only if $x=0$.
  6. 6.
    Our proof is based on that of (BrandenbergKonig13, , Theorem 4.1). First suppose that $A\subseteq\mathbb{R}^{k}$ satisfies
    $\sum_{a\in A}a=0.$ (4)
    For each $a\in A$ we then have $a=\sum_{a^{\prime}\in A\setminus\{a\}}a^{\prime}$ and so
    $\frac{1}{|A|-1}(-a)=\sum_{a^{\prime}\in A\setminus\{a\}}\frac{a^{\prime}}{|A|-1}\in\mathrm{conv}(A\setminus\{a\}).$
    As well, we have for all $a^{\prime}\neq a$ that $a\in A\setminus\{a^{\prime}\}\subseteq\mathrm{conv}(A\setminus\{a^{\prime}\})$. Therefore
    $\frac{(|A|-2)|A|}{|A|-1}a$ $=(|A|-1)a-\frac{1}{|A|-1}a$
    $\in\sum_{a^{\prime}\in A\setminus\{a\}}\mathrm{conv}(A\setminus\{a^{\prime}\})+\mathrm{conv}(A\setminus\{a\})$
    $=\sum_{a^{\prime}\in A}\mathrm{conv}(A\setminus\{a^{\prime}\}).$
    This holds for all $a\in A$, so
    $A\subseteq\frac{|A|-1}{(|A|-2)|A|}\sum_{a^{\prime}\in A}\mathrm{conv}(A\setminus\{a^{\prime}\})=\mathrm{conv}\left(\frac{|A|-1}{(|A|-2)|A|}\sum_{a^{\prime}\in A}(A\setminus\{a^{\prime}\})\right)$
    where we have used the identity $\mathrm{conv}(X+Y)=\mathrm{conv}(X)+\mathrm{conv}(Y)$, see (Schneider14, , Theorem 1.1.2). By part 4. of the Proposition,
    $\delta(A)\leq\mbox{ $\frac{|A|-1}{|A|(|A|-2)}$}\sum_{a\in A}\delta(A\setminus\{a\}).$ (5)
    Now suppose that $\sum_{a\in A}a\neq 0$. Define $a_{0}=\frac{1}{|A|}\sum_{a\in A}a$ and $A_{0}=A-a_{0}$. Then $\sum_{a\in A_{0}}a=0$ and from (5) and translation invariance,
    $\delta(A)$ $=\delta(A_{0})$
     $\leq\mbox{ $\frac{|A_{0}|-1}{|A_{0}|(|A_{0}|-2)}$}\sum_{a\in A_{0}}\delta(A_{0}\setminus\{a\})$
    = $=\mbox{ $\frac{|A|-1}{|A|(|A|-2)}$}\sum_{a\in A}\delta((A-a_{0})\setminus\{a-a_{0}\})$
    = $=\mbox{ $\frac{|A|-1}{|A|(|A|-2)}$}\sum_{a\in A}\delta(A\setminus\{a\})$
    as required.
We now show that the zonotope diversity introduced in Section 2.1 is in fact a sublinear diversity. The length of a zonotope equals its mean width (with the appropriate scaling) so for any two zonotopes $Z_{1},Z_{2}\subseteq\mathbb{R}^{k}$ and $\alpha,\beta\geq 0$ we have that $Z_{1}\subseteq Z_{2}$ implies $\ell(Z_{1})\leq\ell(Z_{2})$ and
$\ell(\alpha Z_{1}+\beta Z_{2})=\alpha\ell(Z_{1})+\beta\ell(Z_{2}).$
The zonotope diversity of finite $A\subseteq\mathbb{R}^{k}$ is
$\delta_{z}(A)=\inf\{\ell(Z):A\subseteq Z,\mbox{ $Z$ a zonotope}\}.$
The induced metric of the zonotope diversity is the Euclidean metric.
@@
Recall that $\delta_{z}(A)$ is the infimum of the lengths of a zonotopes containing $A$. The function $\delta_{z}(A)$ is clearly monotonic, vanishes when $|A|\leq 1$, and is strictly positive when $|A|>1$. Let $\epsilon>0$ be given. Given finite $A$ and $\lambda>0$, let $Z_{A}$ be a zonotope containing $A$ with $\ell(Z_{A})<\delta_{z}(A)+\epsilon$. Then $\lambda Z_{A}$ (with length less than $\lambda\delta_{z}(A)+\lambda\epsilon$) contains $\lambda A$, so $\delta_{z}(\lambda A)\leq\lambda\delta_{z}(A)+\lambda\epsilon$. Letting $\epsilon$ go to zero gives $\delta_{z}(\lambda A)\leq\lambda\delta_{z}(A)$. The other direction ($\lambda\delta_{z}(A)\leq\delta_{z}(\lambda A))$ is obtained similarly and thus gives positive homogeneity. Given finite $A,B\subseteq\mathbb{R}^{k}$ and $\epsilon>0$, let $Z_{A}$ and $Z_{B}$ be zonotopes containing $A$ and $B$ respectively, with $\ell(Z_{A})\leq\delta_{z}(A)+\epsilon/2$ and $\ell(Z_{B})\leq\delta_{z}(B)+\epsilon/2$. Then $Z_{A}+Z_{B}$ is a zonotope containing $A+B$ with length $\ell(Z_{A})+\ell(Z_{B})\leq\delta_{z}(A)+\delta_{z}(B)+\epsilon$. Hence $\delta_{z}(A+B)\leq\delta_{z}(A)+\delta_{z}(B)+\epsilon$. Letting $\epsilon$ go to zero gives sublinearity. By Proposition 1 part 2, $(\mathbb{R}^{k},\delta_{z})$ is a sublinear diversity. ∎
As an example, let $T$ be the three vertices of an equilateral triangle with side length $1$ and centroid at the origin (Fig. 1). Grünbaum (grunbaum1963measures, , pg. 257), citing Kovetz1962SomeExtremalProblems , states that a symmetric convex set containing the equilateral triangle has perimeter at least $2\sqrt{3}$ and that this is bound realized by a hexagon with side length $\frac{1}{\sqrt{3}}$ (Fig. 1). As any symmetric convex polygon in $\mathbb{R}^{2}$ is a zonotope, $\delta_{z}(T)=\delta_{z}(-T)=\sqrt{3}$.
Figure 1: An example illustrating that the zonotope diversity is not linear. $T$ is the set of vertices of an equilateral triangle, centred at the origin. $Z$ is the smallest zonotope containing $T$, which is the same as the smallest zonotope containing T-T. The length of the convex hull of TTT-T is 33, which is smaller than $2\sqrt{3}$, the length of $Z+Z$. Hence $\delta_{z}(T-T)<\delta_{z}(T)+\delta_{z}(-T)$.
The convex hull of TTT-T is the regular hexagon, centered at 0, with side length $1$. This is already a zonotope, so $\delta_{z}(T-T)=3$. We therefore have that $\delta_{z}(T+(-T))<\delta_{z}(T)+\delta_{z}(-T)$, showing that the zonotope diversity is not, in general, linear.
Let $\|\cdot\|$ be a norm on $\mathbb{R}^{k}$ with associated metric $d(x,y)=\|x-y\|$ and unit ball $\mathcal{B}=\{x:\|x\|\leq 1\}$. The Hausdorff distance between two nonempty closed bounded sets $K,L\subset\mathbb{R}^{k}$ can be defined by (Schneider14, , p. 61) :
@@
We extend the function $\delta$ for a sublinear diversity $(X,\delta)$ from finite subsets to bounded $K\subseteq\mathbb{R}^{k}$. Define
@@
    For all finite $A\subseteq\mathbb{R}^{k}$,
@@
    If $(\mathbb{R}^{k},\delta)$ is linear then
    $\delta^{*}(K+L)=\delta^{*}(K)+\delta^{*}(L).$
@@
    Suppose that $K\subseteq[-r,r]^{k}$ for some $r<\infty$. Let $e_{1},\ldots,e_{k}$ be the standard basis for $\mathbb{R}^{k}$ and define
    $V=\sum_{i=1}^{k}\{-re_{i},re_{i}\}$
    so $K\subseteq\mathrm{conv}(V)=[-r,r]^{k}$. By Proposition 1, part 4., $\delta(A)\leq\delta(V)$ for all finite $A\subseteq K$ so
    $\delta^{*}(K)\leq\delta(V)\leq\sum_{i=1}^{k}\delta(\{-re_{i},re_{i}\})<\infty.$
@@
    Let $A$ be a finite subset of $\mathbb{R}^{k}$ and let $K=\mathrm{conv}(A)$. For any finite $A^{\prime}\subseteq K$ we have $\mathrm{conv}(A\cup A^{\prime})=\mathrm{conv}(A)$ so $\delta(A^{\prime})\leq\delta(A\cup A^{\prime})=\delta(A)$ by Proposition 1 (3). Hence
    $\delta(A)\leq\sup\{\delta(A^{\prime}):\mbox{ finite }A^{\prime}\subseteq K\}\leq\delta(A).$
@@
    For any $K$ and $\alpha\geq 0$ we have
    $\delta^{*}(\alpha K)=\sup_{\text{finite }A:\,A\subseteq\alpha K}\delta(A)=\sup_{\text{finite }B:\,\alpha B\subseteq\alpha K}\delta(\alpha B)=\sup_{\text{finite }B:\,B\subseteq K}\alpha\delta(B)=\alpha\delta^{*}(K)$
    where we have used the change of variables $A=\alpha B$.
    Suppose that $(\mathbb{R}^{k},\delta)$ is linear. Given $\epsilon>0$ there are finite $A\subseteq K$ and $B\subseteq L$ such that
    $\delta^{*}(K)<\delta(A)+\epsilon/2\mbox{ and }\delta^{*}(L)<\delta(B)+\epsilon/2.$
    We then have
    $\delta^{*}(K)+\delta^{*}(L)<\delta(A)+\delta(B)+\epsilon=\delta(A+B)+\epsilon\leq\delta^{*}(K+L)+\epsilon$
    as $A+B\subseteq K+L$. Taking $\epsilon\rightarrow 0$ and applying sublinearity gives
    $\delta^{*}(K+L)=\delta^{*}(K)+\delta^{*}(L).$
@@
    By (Schneider14, , Lem. 3.1.1) we have that if $K,L,K\cup L$ and $K\cap L$ are nonempty compact convex subsets then
@@
    Suppose that $K,L$ are bounded nonempty subsets satisfying $d_{H}(K,L)=\lambda$. For any $\epsilon>0$ there is a finite $A\subseteq K$ such that $\delta(A)\leq\delta^{*}(K)<\delta(A)+\epsilon$. We also have $A\subseteq K\subseteq L+\lambda\mathcal{B}$ so there are finite $B\subseteq L$ and $C\subseteq\mathcal{B}$ such that $A\subseteq B+\lambda C$. Hence
@@
Bryant et al. bryant2023diversities also describe an extension of Minkowski diversities from finite sets to bounded sets. They define $\widetilde{\delta}(P)=\delta(\mathrm{vert}(P))$ for any polytope with vertex set $\mathrm{vert}(P)$, and extend that to general bounded convex sets $K$ by defining $\widetilde{\delta}(K)=\lim_{n\rightarrow\infty}\widetilde{\delta}(P_{n})$ for any sequence of polytopes $P_{1},P_{2},\ldots$ converging to $K$ under the Hausdorff metric.
From Proposition 3 part 5. we have that $\widetilde{\delta}$ is well defined for any sublinear diversity. Proposition 3 part 2. gives that $\delta^{*}(P)=\widetilde{\delta}(P)$ for any polytope and from Proposition 3 part 5 we have that $\delta^{*}(K)=\widetilde{\delta}(K)$. Hence $\delta^{*}$ coincides with $\widetilde{\delta}$ for sublinear diversities.
We conclude this section by showing that linear maps of linear (sublinear) semidiversities are linear (resp. sublinear).
Proposition 4.
Let $\phi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ be linear and let $(\mathbb{R}^{n},\delta_{n})$ be a sublinear semidiversity. Then $(\mathbb{R}^{m},\delta_{m})$ given by $\delta_{m}(A)=\delta_{n}(\phi(A))$ is a sublinear semidiversity. If $(\mathbb{R}^{n},\delta_{n})$ is also linear then so is $(\mathbb{R}^{m},\delta_{m})$.
Proof.
As defined, $\delta_{m}$ satisfies (D1) and (D3). For finite $A,B\subseteq\mathbb{R}^{m}$ and $\alpha,\beta\geq 0$ we have
$\delta_{m}(\alpha A+\beta B)=\delta_{n}\big(\phi(\alpha A+\beta B)\big)=\delta_{n}\big(\alpha\phi(A)+\beta\phi(B)\big).$
If $\delta_{n}$ is sublinear then we have
$\delta_{n}\big(\alpha\phi(A)+\beta\phi(B)\big)\leq\alpha\delta_{n}\big(\phi(A)\big)+\beta\delta_{n}\big(\phi(B)\big)=\alpha\delta_{m}(A)+\beta\delta_{m}(B)$
so $(\mathbb{R}^{m},\delta_{m})$ satisfies (D5) and is a sublinear diversity by Proposition 1. If $\delta_{n}$ is also linear then
$\delta_{n}\big(\alpha\phi(A)+\beta\phi(B)\big)=\alpha\delta_{n}\big(\phi(A)\big)+\beta\delta_{n}\big(\phi(B)\big)=\alpha\delta_{m}(A)+\beta\delta_{m}(B)$
and $(\mathbb{R}^{m},\delta_{m})$ is linear. ∎